# Characteristic Guidance: Non-linear Correction for DDPM at Large Guidance Scale

### v1

@@ Abstract
Popular guidance for denoising diffusion probabilistic model (DDPM) linearly combines distinct conditional models together to provide enhanced control over samples. However, this approach overlooks nonlinear effects that become significant when guidance scale is large. To address this issue, we propose characteristic guidance, a novel method that provides non-linear correction for classifier-free guided DDPMs. Such correction forces the guided DDPMs to respect the Fokker-Planck equation of their underlying diffusion process, in a way that is first-principle, training-free, derivative-free, and compatible with existing sampling methods. Experiments show that characteristic guidance is robust to various applications, offers enhanced control over sample generation, suppresses color and exposure issues even for latent space sampling, and can handle physics problems such as the phase transitions.
@@ 1 Introduction
The diffusion model (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b) is a family of generative models that produce high-quality samples by mimicking the diffusion processes of molecules. Denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) is one of the most popular diffusion models whose diffusion process can be viewed as a sequence of denoising steps. It is the core of several large text-to-image models that have been widely adopted in AI art production, such as the latent diffusion model (stable diffusion) (Rombach et al., 2021) and DALL·E(Betker et al., 2023). However, providing precise control over these models to achieve the desired content and style, which is essential for successful AI art production, remains a challenging problem.
DDPMs control the generation of samples by learning conditional distributions that are specified by sample contexts. However, such conditional DDPMs has weak control strength hence do not work well to generate desired content (Luo, 2022). Guidance is a technique providing adjustable control strength at the cost of sample diversity. The two most popular guidance are the classifier guidance (Song et al., 2020b; Dhariwal & Nichol, 2021) and the classifier-free guidance (Ho, 2022). Classifier guidance requires an additional classifier to guide the sampling process, which is difficult to implement on non-classification tasks such as text-to-image generation. Classifier-free guidance avoided the use of classifier by combining conditional and unconditional DDPM linearly together, weighted by a guidance scale parameter that determine the control strength. However, classifier-free guidance produces highly saturated and unnatural images (Ho, 2022; Saharia et al., 2022) when the control strength is large.
Our work attribute classifier-free guidance’s irregular behavior at large control strength to its theoretical flaw: The neglect of the non-linearity of DDPM. Theoretically, both the conditional and unconditional DDPM satisfies the score Fokker-Planck equaiton (Lai et al., 2022) but linear combination of them does not. To address this issue, we propose characteristic guidance as a non-linear corrected guidance technique. It utilizes the method of characteristics to solve the score Fokker-Planck equation approximately, correcting the classifier-free guidance in a first-principle, training-free, derivative-free way that is compatible with existing sampling methods. Our experiments show that characteristic guidance is robust to various applications, offers enhanced control over sample generation, suppresses color and exposure issues even for latent space sampling, and can handle physics problems such as phase transitions.
Figure 1: Comparative visualization of volcano images generated from latent diffusion model using Classifier Free Guidance (CF) versus Characteristic Guidance (CH) without cherry-picking. With increasing guidance scale ($\omega$), CF-guided images exhibit severe color cast and underexposure issues. In contrast, CH guidance produces images with consistent color and exposure. Furthermore, the CH guidance progressively enhances volcanic features such as smoke and lava, while CF guidance remains less responsive to these details.
@@ 2 Related Works
Diffusion Models and Guidance Techniques
: Diffusion models, such as the widely used Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), have emerged as a powerful class of generative models to create high-quality samples (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b). Controlling these models, especially in AI art production, is a critical challenge. Various methods have been proposed for guiding diffusion models, including approaches utilizing classifiers (Song et al., 2020b; Dhariwal & Nichol, 2021) and classifier-free strategies (Ho, 2022). Additionally, recent advancements have focused on enhancing control styles (Bansal et al., 2023; Yu et al., 2023) and improving sampling quality (Hong et al., 2022; Kim et al., 2022). Most existing techniques employ guidance as a linear combination of scores or equivalent noises, which hurts sampling quality at large guidance scales. Our work seeks to tackle this issue by addressing the non-linear effects in DDPMs in a training-free way.
Fokker-Planck Equation and DDPMs
The Fokker-Planck equation, central to the diffusion process of DDPMs (Song et al., 2020b), is approximately adhered to by most trained DDPMs. Enhancing DDPM performance by promoting greater alignment with the Fokker-Planck equation has shown promise (Lai et al., 2022), as has maintaining the conservation property of scores (Chao et al., 2022). Yet, leveraging insights from the Fokker-Planck equation typically requires computationally intensive calculations of derivatives of the DDPM model. Additionally, aligning guided DDPMs with the Fokker-Planck equation remains an unresolved challenge. Our research seeks to correct guided DDPMs in a derivative-free approach, harnessing the Fokker-Planck equation’s benefits while minimizing computational demands.
Fast Sampling Methods
The inherent slow sampling speed of DDPMs, necessitating hundreds to thousands of model evaluations per sample, is a significant limitation. To mitigate this, various strategies have been developed to expedite the sampling process. These include approximating the diffusion process (Song et al., 2020b; Liu et al., 2022; Song et al., 2020a; Zhao et al., 2023) and employing advanced ODE/PDE solvers like Runge–Kutta and predictor-corrector methods (Karras et al., 2022; Lu et al., 2022; Zhao et al., 2023). All these methods relies on scores or predicted noises to do the de-noising steps. Our goal is to correct guided DDPMs by adjusting scores or predicted noises in a way that is compatible with these fast sampling technique.
@@ 3 Background
3.1 The Denoising Diffusion Probabilistic Model (DDPM)
DDPM models distributions of images $p(\mathbf{x})$ by recovering an original image $\mathbf{x}_{0}\sim p(\mathbf{x})$ from one of its contaminated versions $\mathbf{x}_{i}$. The contaminated image is a linear combination of the original image and a gaussian noise
@@ 3.1 The Denoising Diffusion Probabilistic Model (DDPM)
where $n$ is the total diffusion steps, $\bar{\alpha}_{i}$ is the contamination weight at a time $t_{i}\in[0,T]$ in the forward diffusion process, and $\bar{\boldsymbol{\epsilon}}_{i}$ is a standard Gaussian random noise to be removed. More detailed information could be found in Appendix A.
@@
The denoising neural network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})$ is later used in the backward diffusion process to generate image samples from pure Gaussian noises by removing noises step by step.
@@ 3.2 Conditional DDPM and Classifier-Free Guidance
Conditional DDPMs, which generate images based on a given condition $\mathbf{c}$, model the conditional image distribution $p(\mathbf{x}|\mathbf{c})$. One can introduce the dependency on conditions by extending the denoising neural network to include the condition $\mathbf{c}$, represented as $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})$.
@@
Classifier-free guidance is a technique for conditional image generation that trades off control strength and image diversity. It aims to sample from the distribution (Ho, 2022):
@@
where $\omega>0$ is the guidance scale. When $\omega$ is large, the guided DDPM produces samples that have the highest condition likelihood $p(\mathbf{c}|\mathbf{x})$.
@@
The connection between scores and the denoising neural network (Appendix B) inspires the classifier free guidance to use the following guided denoising neural network $\boldsymbol{\epsilon}_{CF}$:
The classifier free guidance exactly computes the de-noising neural network of $p(\mathbf{x}|\mathbf{c},\omega)$ at time $t_{0}=0$ (Appendix B). However, it is not a good approximation of the theretical DDPM modeling $p(\mathbf{x}|\mathbf{c},\omega)$ for most of diffusion time steps $t_{i}>0$, especially when the guidance scale $\omega$ is large. This is because networks $\boldsymbol{\epsilon}$ respect the Fokker-Plank equation of the diffusion process (Appendix C) in a non-linear way hence cannot be linearly combined easily.
4 Methogology
@@ 4.1 The Characteristic Guidance
Our work aims to provide non-linear corrections to the classifier-free guidance, making it approximately satisfy the Fokker-Planck equation in a training-free and derivative-free way. We propose the characteristic guidance as non-linear corrected classifier-free guidance:
@@
We determine the correction $\Delta\mathbf{x}$ from the following non-linear relation:
$\begin{split}\Delta\mathbf{x}&=\mathbf{P}\circ\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{2},t_{i})-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{1}|\mathbf{c},t_{i})\right)\sigma_{i},\end{split}$ (6)
where $\sigma_{i}=\sqrt{1-\bar{\alpha}_{i}}$ is a scale parameter and the operator $\mathbf{P}$ could be an arbitrary orthogonal projection operator. In practice, the equation (6) could be solved with fix-point iteration methods (Appendix D).
The operator $\mathbf{P}$ should be theoretically the identity, but we found that orthogonal projections that serve as regularization greatly accelerates the computation when sampling natural images. For pixel space diffusion model, we suggest the operator $\mathbf{P}$ to be projection from the input to its channel-wise mean. For latent space diffusion model, we suggest the operator $\mathbf{P}$ to be the channel-wise projection to the residual vector $\mathbf{g}=\Delta\mathbf{x}-\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})\right)\sigma_{i}$. For low dimensional cases that are not images, we suggest the operator $\mathbf{P}$ to be identity.

### v5

# Characteristic Guidance: Non-linear Correction for Diffusion Model at Large Guidance Scale

@@ Abstract
Popular guidance for denoising diffusion probabilistic model (DDPM) linearly combines distinct conditional models together to provide enhanced control over samples. However, this approach overlooks nonlinear effects that become significant when guidance scale is large. To address this issue, we propose characteristic guidance, a guidance method that provides first-principle non-linear correction for classifier-free guidance. Such correction forces the guided DDPMs to respect the Fokker-Planck (FP) equation of diffusion process, in a way that is training-free and compatible with existing sampling methods. Experiments show that characteristic guidance enhances semantic characteristics of prompts and mitigate irregularities in image generation, proving effective in diverse applications ranging from simulating magnet phase transitions to latent space sampling.
@@ 1 Introduction
Figure 1: Comparative visualization of images sampled from Stable diffusion XL (Podell et al., 2023) between Classifier Free Guidance and Characteristic Guidance (Model name: animagineXL 3.0 (CagliostroLab, 2023), Seeds: 0,1,2,3). By addressing the non-linear effects of the FP equation, characteristic guidance demonstrates ability to mitigate irregularity in color, exposure and anatomy and enhancing prompt’s semantic characteristics (e.g., “grass” depicted in the right-hand example).
@@
The diffusion model (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b) is a family of generative models that produce high-quality samples utilizing the diffusion processes of molecules. Denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) is one of the most popular diffusion models whose diffusion process can be viewed as a sequence of denoising steps. It is the core of several large text-to-image models that have been widely adopted in AI art production, such as the latent diffusion model (stable diffusion) (Rombach et al., 2021) and DALL·E (Betker et al., 2023). DDPMs control the generation of samples by learning conditional distributions that are specified by sample contexts. However, such conditional DDPMs has weak control over context hence do not work well to generate desired content (Luo, 2022).
Guidance techniques, notably classifier guidance (Song et al., 2020b; Dhariwal & Nichol, 2021) and classifier-free guidance (Ho, 2022), provide enhanced control at the cost of sample diversity. Classifier guidance, requiring an additional classifier, faces implementation challenges in non-classification tasks like text-to-image generation. Classifier-free guidance circumvents this by linearly combining conditional and unconditional DDPM weighted by a guidance scale parameter. However, large guidance scale often leads to overly saturated and unnatural images (Ho, 2022). Though techniques like dynamic thresholding (Saharia et al., 2022) can handle color issues by clipping out-of-range pixel values, a systematic solution for general sampling tasks, including latent space diffusion (Rombach et al., 2021) and those beyond image generation, remains absent.
This paper proposes characteristic guidance as a systematic solution to the large guidance scale issue by addressing the non-linearity neglicted in classifier-free guidance. The name characteristic serves a dual purpose: Technically, characteristic guidance utilizes the method of characteristics in solving the score Fokker-Planck (FP) equation (Lai et al., 2022) for diffusion process, addressing its non-linearity to which the classifier-free guidance fails to adhere. It demonstrates clear theoretical advantages on Gaussian models in removing large guidance scale irregularities. Experimentally, characteristic guidance notably enhances the semantic characteristics of images, ensuring closer alignment with specified conditions. This is evident in our image generation experiments on CIFAR-10, ImageNet 256, and Stable diffusion (Fig.1). Notably, it also mitigates color and exposure issues common at large-scale guidance. Furthermore, characteristic guidance is compatible with any continuous data type and is robust across continuous diverse applications, ranging from simulating magnet phase transitions to latent space sampling.
@@ 2 Related Works
2.1 Diffusion Models and Guidance Techniques
Diffusion models, notably Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), have become pivotal in generating high-quality AI art (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b). Various guidance methods, including classifier-based (Song et al., 2020b; Dhariwal & Nichol, 2021) and classifier-free techniques (Ho, 2022), have been developed. Recent efforts aim to refine control styles (Bansal et al., 2023; Yu et al., 2023) and enhance sampling quality (Hong et al., 2022; Kim et al., 2022). However, current guidance methods suffer from quality degradation and color saturation issues at high guidance scales. Dynamic thresholding (Saharia et al., 2022) attempts to address this by normalizing quantiles of pixels, but its effectiveness is limited to color-based image pixels, falling short in general tasks like latent space sampling. To suppress artifacts, community later had to use ”scale mimic” that mixes dynamic thresholded latents with low guidance scale latents (Mcmonkeyprojects, 2024). Instead of these workarounds, our research aims to address quality issues of classifier-free guidance theoretically and systematically for general tasks, by tackling the non-linear aspects of DDPMs.
2.2 Fast Sampling Methods
The significant challenge posed by the slow sampling speed of DDPMs, requiring numerous model evaluations, has spurred the development of accelerated sampling strategies. These strategies range from diffusion process approximations (Song et al., 2020b; Liu et al., 2022; Song et al., 2020a; Zhao et al., 2023) to employing advanced solvers like Runge–Kutta and predictor-corrector methods (Karras et al., 2022; Lu et al., 2022; Zhao et al., 2023). These methods primarily utilize scores or predicted noises for de-noising. Our research focuses on designing a high guidance scale correction that seamlessly integrates with these fast sampling methods, primarily through the modification of scores or predicted noises.
@@ 3 Background
3.1 DDPM and the FP Equation
DDPM models distributions of images $p(\mathbf{x})$ by recovering an original image $\mathbf{x}_{0}\sim p(\mathbf{x})$ from one of its noise contaminated versions $\mathbf{x}_{i}$. The noise contaminated image is a linear combination of the original image and a gaussian noise
@@ 3.1 DDPM and the FP Equation
where $n$ is the total diffusion steps, $\bar{\alpha}_{i}$ is the contamination weight at a time $t_{i}\in[0,T]$ in the forward diffusion process, and $\bar{\boldsymbol{\epsilon}}_{i}$ is a standard Gaussian random noise. More detailed information can be found in Appendix A.
@@
This neural network is later used in the backward diffusion process to generate image samples from pure Gaussian noises. In the ideal scenario where diffusion time steps among $t_{i}$ are infinitesimal, the optimal solution for (2) is defined by:
$\boldsymbol{\epsilon}(\mathbf{x},t)=\operatorname*{argmin}_{\boldsymbol{\epsilon}_{\theta}}L(\boldsymbol{\epsilon}_{\theta}).$ (3)
The forward process (1) places stringent constraints on its permissible forms. First, $\boldsymbol{\epsilon}(\mathbf{x},t_{i})$ is proportional to the score function $\mathbf{s}(\mathbf{x},t_{i})=\nabla_{\mathbf{x}_{i}}\log p(\mathbf{x}_{i})$ (Yang et al., 2022) (Appendix A). Second, the score function is a solution of the score FP equation (Lai et al., 2022), which could be rewritten (Appendix C) in terms of $\boldsymbol{\epsilon}(\mathbf{x},t)$ as
$\frac{\partial\boldsymbol{\epsilon}}{\partial t}=\frac{1}{2}\left(\mathcal{L}\boldsymbol{\epsilon}-\frac{1}{\sigma(t)}\nabla_{\mathbf{x}}\|\boldsymbol{\epsilon}\|^{2}_{2}\right),$ (4)
in which $\mathcal{L}\boldsymbol{\epsilon}=\nabla_{\mathbf{x}}(\boldsymbol{\epsilon}\cdot\mathbf{x})+\nabla_{\mathbf{x}}^{2}\boldsymbol{\epsilon}+\frac{1-\sigma(t)^{2}}{\sigma(t)^{2}}\boldsymbol{\epsilon}$ and $\sigma(t)=\sqrt{1-e^{-t}}$. These constraints lay the foundation for the duality between forward and backward diffusion processes, which is essential for successful sampling from $p(\mathbf{x})$.
@@ 3.2 Conditional DDPM and Classifier-Free Guidance
Conditional DDPMs, which generate images based on a given condition $\mathbf{c}$, model the conditional distribution $p(\mathbf{x}|\mathbf{c})$ with a denoising neural network represented as $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})$. However, the training data in practice might only have weak or noised information about condition $\mathbf{c}$, therefore we need a way to enhance the control strength in this situation.
@@
Guidance (Song et al., 2020b; Dhariwal & Nichol, 2021; Ho, 2022) is a technique for conditional image generation that trades off control strength and image diversity. It generally aims to sample from the distribution:
$p(\mathbf{x}|\mathbf{c},\omega)\propto p(\mathbf{x}|\mathbf{c})^{1+\omega}p(\mathbf{x})^{-\omega}\propto p(\mathbf{c}|\mathbf{x})^{1+\omega}p(\mathbf{x}),$ (5)
@@
where $\omega>0$ is the guidance scale. When $\omega$ is large, this distribution concentrates on samples that have the highest conditional likelihood $p(\mathbf{c}|\mathbf{x})$.
@@
Classifier free guidance (Ho, 2022) use the following guided denoising neural network $\boldsymbol{\epsilon}_{CF}$, deduced from (5) using $\boldsymbol{\epsilon}\propto\nabla\log p$, to approximately sample from $p(\mathbf{x}|\mathbf{c},\omega)$:
$\boldsymbol{\epsilon}_{CF}(\mathbf{x}|\mathbf{c},t_{i},\omega)=(1+\omega)\ \boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})-\omega\ \boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i}).$ (6)
It exactly computes $\boldsymbol{\epsilon}(\mathbf{x}|\mathbf{c},t_{i},\omega)$, the de-noising neural network of $p(\mathbf{x}|\mathbf{c},\omega)$, at time $t_{i}=0$ (Appendix B). However, it is not a good approximation for $t_{i}>0$ as it deviates from the FP equation (4) when $\omega$ is large.
3.3 Guidance’s Deviation from the FP Equation
We measure the deviation of a guidance method from the FP equation using the mixing error:
$\mathbf{e}_{m}(\boldsymbol{\epsilon},\mathbf{x},t)=\frac{\partial\boldsymbol{\epsilon}}{\partial t}-\frac{1}{2}\left(\mathcal{L}\boldsymbol{\epsilon}-\frac{\nabla_{\mathbf{x}}\|\boldsymbol{\epsilon}\|^{2}_{2}}{\sigma(t)}\right).$ (7)
in which $\boldsymbol{\epsilon}$ is the guided denoising neural network under infinitesimal diffusion time steps.
For classifier-free guidance, the mixing error comes from the non-linear term of (4) as
$\begin{split}&\mathbf{e}_{m}(\boldsymbol{\epsilon}_{CF},\mathbf{x},t)=\frac{1}{\sigma(t)}\nabla_{\mathbf{x}}(\|\boldsymbol{\epsilon}_{CF}(\mathbf{x}|\mathbf{c},t_{i},\omega)\|^{2}_{2}-\\ &(1+\omega)\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})\|^{2}_{2}+\omega\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})\|^{2}_{2}),\end{split}$ (8)
whose amplitude escalates with an increase in $\omega$ and further intensifies as $\sigma(t)$ decreases.
3.4 The Method of Characteristics
The method of characteristics (Evans, 2022) provides an analytical approach to solving certain kind of non-linear partial differential equations (PDE). It employs characteristic lines, a family of time-dependent trajectories defined by $\mathbf{x}_{t}=\mathbf{x}(t)$, along which the equation
$\boldsymbol{\epsilon}(\mathbf{x}_{t},t)=\mathbf{F}_{t}(\boldsymbol{\epsilon}(\mathbf{x}_{0},0))$ (9)
holds, where $\boldsymbol{\epsilon}$ is a solution of the PDE and $\mathbf{F}_{t}$ is a deterministic and invertible function. This implies that the value of $\boldsymbol{\epsilon}(\mathbf{x}_{t},t)$ can be inferred from its initial condition $\boldsymbol{\epsilon}(\mathbf{x}_{0},0)$ at a specific point $\mathbf{x}_{0}$, determined by the characteristic line itself.
The method of characteristics is applicable to (4) under certain assumptions, whose characteristic lines are detailed in (50) of Appendix D. Particularly, the function $\mathbf{F}_{t}$ for (4) is a linear function.
4 Problem Description
Our method emphasizes the importance of adhering to the Fokker-Planck (FP) equation for two main reasons: Theoretically, the deviation from the FP equation leads to mixing errors, disrupting the equivalence between forward and backward diffusion processes and resulting in sample generation irregularities. Empirically, as we will shown in Sec.6.1, aligning with the FP equation helps to remove irregularities of classifier-free guidance.
We focus on a guided DDPM that generates samples from $p(\mathbf{x}|\mathbf{c},\omega)$ (5), constructing its denoising neural network $\boldsymbol{\epsilon}_{CH}(\mathbf{x}|\mathbf{c},t_{i},\omega)$ from two known networks: $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})$ and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})$. The desired properties of $\boldsymbol{\epsilon}_{CH}$ include:
  • $\boldsymbol{\epsilon}_{CH}$ has no mixing error (satisfying FP equation);
  • $\boldsymbol{\epsilon}_{CH}$ requires no training.
Constructing such $\boldsymbol{\epsilon}_{CH}$ is highly non-trivial as it requires tackling the non-linear term in the Fokker-Planck equation.
5 Methodology
@@ 5.1 The Characteristic Guidance
We propose the characteristic guidance as non-linear corrected classifier-free guidance:
@@
The correction $\Delta\mathbf{x}$ is determined from the training-free non-linear relation:
$\Delta\mathbf{x}=\mathbf{P}\circ\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{2},t_{i})-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{1}|\mathbf{c},t_{i})\right)\sigma_{i},$ (11)
where $\sigma_{i}=\sqrt{1-\bar{\alpha}_{i}}$ is a scale parameter and the operator $\mathbf{P}$ could be an arbitrary orthogonal projection operator.
Equation (11) can be solved using the fixed-point iteration method (Evans, 2022) (Appendix E). For practical efficiency, we propose accelerated fixed-point iteration algorithms, Alg.2 and Alg.3, to minimize the iterations needed for convergence.
The projection operator $\mathbf{P}$ should be theoretically the identity, but we found that orthogonal projections that serve as regularization greatly accelerates the computation. In practice, $\mathbf{P}$ acts channel-wisely and is specified by a vector $\mathbf{g}$:
$\mathbf{P}_{\mathbf{g}}\circ\mathbf{v}=\frac{\mathbf{g}\cdot\mathbf{v}}{\mathbf{g}\cdot\mathbf{g}}\mathbf{g}$ (12)
For pixel space diffusion model, we suggest the operator $\mathbf{g}=\mathbf{1}$ as projection to the channel-wise mean. For latent space diffusion model, we suggest $\mathbf{g}=\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})\right)\sigma_{i}$. For low dimensional cases that are not images, we suggest the operator $\mathbf{P}$ to be identity.